OpenStudy (anonymous):

if y=x^2+3/x, find x as a function of y

OpenStudy (shubhamsrg):

its a cubic equation hmm you comfortable with those?

OpenStudy (shubhamsrg):

oh wait, is it (x^2 +3)/x ?

OpenStudy (anonymous):

no its y=x^2+3 over x or y=x^2+3/3 find x as the function of y